Please help me solve:
(b-4)^2=16
I know I have to use the quadratic formula, but I can't figure it out

Answer by stanbon(75887)

Please help me solve:
(b-4)^2=16
----
b-4 = sqrt(16) or b-4 = -sqrt(16)
b = 4+4 or b = 4-4
---
b = 8 or b = 0
===================
Cheers,
Stan H.